Piqua Ohio is known for its numerous stoplights. As you drive through downtown Piqua you notice there are 6 stoplights. Each stoplight works independently of each other and each one has a 30% change of being red. What is the probability of hitting less than 3 red lights?
Answer #1

Given that ,

p = 0.30

1 - p = 1 - 0.30= 0.70

n = 6

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X < 3) = P(X=0) +P(X=1)+P(X=2)

= (6 C 0) * 0.300 * (0.70)6+(6 C 1) * 0.301 * (0.70)5+(6 C 2) * 0.302 * (0.70)4

probability =0.7443